# Interpretable Hybrid-Rule Temporal Point Processes

### v1

@@ Abstract
Temporal Point Processes (TPPs) are widely used for modeling event sequences in various medical domains, such as disease onset prediction, progression analysis, and clinical decision support. Although TPPs effectively capture temporal dynamics, their lack of interpretability remains a critical challenge. Recent advancements have introduced interpretable TPPs. However, these methods fail to incorporate numerical features, thereby limiting their ability to generate precise predictions. To address this issue, we propose Hybrid-Rule Temporal Point Processes (HRTPP), a novel framework that integrates temporal logic rules with numerical features, improving both interpretability and predictive accuracy in event modeling. HRTPP comprises three key components: basic intensity for intrinsic event likelihood, rule-based intensity for structured temporal dependencies, and numerical feature intensity for dynamic probability modulation. To effectively discover valid rules, we introduce a two-phase rule mining strategy with Bayesian optimization. To evaluate our method, we establish a multi-criteria assessment framework, incorporating rule validity, model fitting, and temporal predictive accuracy. Experimental results on real-world medical datasets demonstrate that HRTPP outperforms state-of-the-art interpretable TPPs in terms of predictive performance and clinical interpretability. In case studies, the rules extracted by HRTPP explain the disease progression, offering valuable contributions to medical diagnosis. The code is available at https://anonymous.4open.science/r/HRTPP-0315.
@@ II-2 Rule Mining
II-A Temporal Point Processes
@@
II-B Temporal Logic Rules
Temporal logic rules consist of two parts: temporal predicates and relations. Define a set of predicates as $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n},Y\}$, where each predicates $X_{u}$ is an independent variable, $Y$ is a target variable. Define a set of temporal relations as $\mathcal{C}$, which define how the variables are related over time. Logic rules are in the form of ”if-then”. A generalized temporal logic rule $R$ can be expressed in the form:
@@ III Methodology
III-A Rule-Based Feature Construction
@@
III-A1 Definitions of Temporal Rules
@@ III-A2 Rule Encoder
III-A3 Numerical Feature Encoder
@@
III-B Hybrid-Rule Temporal Point Processes
@@ III-B1 Basic Intensity
III-B2 Rule-Based Intensity
@@
III-B3 Numerical Feature Intensity
@@
$\begin{split}\lambda(t|\mathcal{H}_{t})&=\text{Softplus}\left(\lambda_{\text{base}}(t)+\lambda_{\text{rule}}(t)+\lambda_{\text{num}}(t)\right)\\ &=\text{Softplus}\Big{(}\lambda_{0}+\sum_{R_{j}\in\mathcal{R}}\alpha_{j}\sum_{t_{j}\in\mathcal{T}j}d{{}_{\text{rule}}}(t-t_{j})\\ &\quad+\sum_{k=1}^{K}\beta_{k}m_{k}\sum_{i:k_{i}=k}v_{i}d_{\text{num}}(t-t_{i})\Big{)}.\end{split}$ (8)
@@ III-C2 Event Prediction
III-D Rule Mining and Optimization
@@
III-D1 Rule Candidate Generation
@@
III-D2 Rule Optimization via Bayesian Optimization
@@
IV Experiments
IV-A Experimental Setup
@@ IV-C Rule Analysis
IV-C1 Rule Expression Analysis
@@
IV-C2 Rule Distribution Analysis
@@ IV-E Case Study
Figure 4: Intensity over time for a CAD patient.

### v2

@@ Abstract
Temporal Point Processes (TPPs) are widely used for modeling event sequences in various medical domains, such as disease onset prediction, progression analysis, and clinical decision support. Although TPPs effectively capture temporal dynamics, their lack of interpretability remains a critical challenge. Recent advancements have introduced interpretable TPPs. However, these methods fail to incorporate numerical features, thereby limiting their ability to generate precise predictions. To address this issue, we propose Hybrid-Rule Temporal Point Processes (HRTPP), a novel framework that integrates temporal logic rules with numerical features, improving both interpretability and predictive accuracy in event modeling. HRTPP comprises three key components: basic intensity for intrinsic event likelihood, rule-based intensity for structured temporal dependencies, and numerical feature intensity for dynamic probability modulation. To effectively discover valid rules, we introduce a two-phase rule mining strategy with Bayesian optimization. To evaluate our method, we establish a multi-criteria assessment framework, incorporating rule validity, model fitting, and temporal predictive accuracy. Experimental results on real-world medical datasets demonstrate that HRTPP outperforms state-of-the-art interpretable TPPs in terms of predictive performance and clinical interpretability. In case studies, the rules extracted by HRTPP explain the disease progression, offering valuable contributions to medical diagnosis. The code is available at https://github.com/yy-c/HRTPP.
@@ II-2 Rule Mining
III Background
TPPs are widely used to interpret complex event sequences over time. For example, in ICU monitoring, physicians analyze the progression of symptoms and lab results to make decisions. Understanding the logic rules among medical events is crucial for accurate diagnosis and treatment. This section briefly reviews temporal point processes and temporal logic rules.
III-A Temporal Point Processes
@@
III-B Temporal Logic Rules
Temporal logic rules consist of two parts: temporal predicates and relations. Define a set of temporal predicates as $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n},Y\}$, where each predicates $X_{u}$ is an independent variable, $Y$ is a target variable. Define a set of temporal relations as $\mathcal{C}$, which define how the variables are related over time. Logic rules are in the form of ”if-then”. A generalized temporal logic rule $R$ can be expressed in the form:
@@ IV Methodology
IV-A Rule-Based Feature Construction
@@
IV-A1 Definitions of Temporal Rules
@@ IV-A2 Rule Encoder
IV-A3 Numerical Feature Encoder
@@
IV-B Hybrid-Rule Temporal Point Processes
@@ IV-B1 Basic Intensity
IV-B2 Rule-Based Intensity
@@
IV-B3 Numerical Feature Intensity
@@
$\begin{split}\lambda(t|\mathcal{H}_{t})&=\text{Softplus}\left(\lambda_{\text{base}}(t)+\lambda_{\text{rule}}(t)+\lambda_{\text{num}}(t)\right)\\ &=\text{Softplus}\left(\lambda_{0}+\sum_{R_{j}\in\mathcal{R}}\alpha_{j}\sum_{t_{j}\in\mathcal{T}j}d{{}_{\text{rule}}}(t-t_{j})+\sum_{k=1}^{K}\beta_{k}m_{k}\sum_{i:k_{i}=k}v_{i}d_{\text{num}}(t-t_{i})\right).\end{split}$ (8)
@@ IV-C2 Event Prediction
IV-D Rule Mining and Optimization
@@
IV-D1 Rule Candidate Generation
@@
IV-D2 Rule Optimization via Bayesian Optimization
@@
V Experiments
V-A Experimental Setup
@@ V-C Rule Analysis
V-C1 Rule Expression Analysis
@@
V-C2 Rule Distribution Analysis
@@ V-E Case Study
Figure 4: Intensity over time for a CAD patient. The red curve represents the intensity of the target variable, reflecting the probability of death. Key events affecting the mortality are annotated below the curve, while the timing of logic rules is marked above. The entire figure illustrates the patient’s full clinical progression in the ICU.